what is the solution of finding the range and the domain in these equations?
1. (x,y)/y=7x-1
2. (x,y)/y=-4x-8
3. y=-8
4. x=1/4
5. (x,y)/y=4x raise to 2
6. ( x,y)/y=-5xsquared+7
7. (x,y)/x=ysquared+3
i hope you'll answer it immediately because i need it now. thank you.

Answer by stanbon(75887)

what is the solution of finding the range and the domain in these equations?
1. (x,y)/y=7x-1
Domain is all Real numbers; Range is all Real numbers
2. (x,y)/y=-4x-8
Domain and Range are all Real numbers
3. y=-8
Domain is all Real numbers; Range is {-8}
4. x=1/4
Domain is {1/4}; Range is all Real numbers
5. (x,y)/y=4x raise to 2
Domain is all Real numbers: Range is all Real >=0
6. ( x,y)/y=-5xsquared+7
Domain is all Real numbers : Range is all Real <=0
7. (x,y)/x=ysquared+3
Domain is all Real numbers >=3; Range is all Real numbers
========
Cheers,
Stan H.